# Beyond Patches: Mining Interpretable Part-Prototypes for Explainable AI

### v1

@@ A.1. Implementation Details
A.2. Additional Ablation Studies
A.2.1. Effect of Hyper-parameters
@@
Figure 8. Impact of hyper-parameters $\alpha$ and $\beta$ on classification accuracy (C Acc) and faithfulness score F(3). Our method maintains strong performance across a wide range of values, showing robustness and stability in both predictive accuracy and interpretability.
@@
Figure 9. The model’s classification accuracy in different values of $m1$ and $m2$.
Figure 10. Visualization of activated concepts in test data and showing some examples in the training set for StanfordCars datasets. Each concept is shown with an index and percentage of explanation for the model decision.
Figure 11. Visualization of activated concepts in test data and showing some examples in the training set for CUB-200 Birds datasets. Each concept is shown with an index and percentage of explanation for the model decision.
Figure 12. Visualization of non-prototypical activated concepts in test data and showing some examples in the training set for (left) Stanford Cars and (right) CUB-200 Birds datasets. Similar to ProtoPNet and PIPNet, we select the patch for the region of the activated concept.
A.2.2. Effect of Non-Prototypical Concepts
To evaluate the impact of different concept vector types on model performance and interpretability, we conduct an ablation study using prototypical ($\mathbf{z^{i}}$) and non-prototypical ($\mathbf{g^{i}}$) concept vectors. As shown in Table 4, using only non-prototypical vectors significantly reduces both classification accuracy and faithfulness scores. In contrast, employing only prototypical vectors retains higher classification performance and strong concept-based faithfulness. The full PCMNet, which combines both concept types, achieves the best results across both datasets, suggesting that the interplay between prototypical and non-prototypical representations enhances both prediction accuracy and interpretability.
@@ C Acc F(1)-F(5) C Acc F(1)-F(5)
Table 4. Impact of prototypical ($\mathbf{z^{i}}$) and non-prototypical ($\mathbf{g^{i}}$) concept vectors on PCMNet’s classification accuracy (C Acc) and faithfulness scores (F(1)-F(5)) on Cars and Birds datasets. The full model achieves superior performance, highlighting the benefit of combining both concept types.
A.3. Additional Qualitative Results
More results of the PCMNet prototypical concept are shown in Figs. 10 and 11. 12 illustrates examples of non-prototypical concepts discovered by our model for the Cars (left) and CUBs (right) datasets. For each dataset, the concept location on a test image is highlighted with a green bounding box, while corresponding activations in training examples are shown on the right using red bounding boxes. These concepts provide complementary cues to prototypical ones. For example, in the first row of Cars, the ”cargo” attributes are not well represented in prototypical concepts since they appear in a small subset of vehicles, yet they are highly discriminative for decision-making. Similarly, the last row in CUBs shows the ”webbed feet” concept, which is especially relevant and explainable for Albatross species, while foot prototypes are not sufficiently distinctive across the dataset. These findings highlight the value of incorporating non-prototypical concepts in improving both interpretability and robustness of the model.

### v2

@@ Appendix A Supplementary: Additional Experiment and Results
Appendix B: Additional Ablation Studies
Number of Parts ($K$)
We assess the impact of the number of parts on classification accuracy, faithfulness, and size of concept prototypes in Table 4. As the model extracts more prototypical concepts, its decisions become more aligned with the underlying features, enhancing interpretability. Interestingly, when the number of prototypes is smaller, the faithfulness scores for the top concepts tend to be higher. This is because the model relies on fewer, more compact concepts ($d_{c}$), concentrating its decisions on a smaller number of more influential ones.
K Cars Birds
C Acc F(1)-F(5) $d_{c}$ C Acc F(1)-F(5) $d_{c}$
0 84.1 1.54 - 6.61 - 83.2 1.21-5.56 -
4 83.5 79.96 - 91.01 1703 83.5 71.42-85.64 1786
5 84.1 70.11 – 91.33 1979 84.9 66.7 - 85.9 2018
6 85.2 66.02 – 90.65 2345 85.7 59.9 - 86.7 2597
7 86.7 62.47 – 91.25 2767 86.0 55.7 - 87.3 2954
8 87.5 59.33 – 91.73 3092 85.1 53.0 - 86.4 3242
9 86.4 55.10 – 91.52 3277 84.4 52.5-86.1 3470
Table 4: Impact on performance of the number of prototypes.
Effect of Hyper-parameters.
@@ C Acc F(1)-F(5) $d_{c}$ C Acc F(1)-F(5) $d_{c}$
Figure 7: Impact of hyper-parameters $\alpha$ and $\beta$ on classification accuracy (C Acc) and faithfulness score F(3). Our method maintains strong performance across a wide range of values, showing robustness and stability in both predictive accuracy and interpretability.
@@
Figure 8: The model’s classification accuracy in different values of $m1$ and $m2$.
Effect of Non-Prototypical Concepts
To evaluate the impact of different concept vector types on model performance and interpretability, we conduct an ablation study using prototypical ($\mathbf{z^{i}}$) and non-prototypical ($\mathbf{g^{i}}$) concept vectors. As shown in Table 5, using only non-prototypical vectors significantly reduces both classification accuracy and faithfulness scores. In contrast, employing only prototypical vectors retains higher classification performance and strong concept-based faithfulness. The full PCMNet, which combines both concept types, achieves the best results across both datasets, suggesting that the interplay between prototypical and non-prototypical representations enhances both prediction accuracy and interpretability. In other words, the vector $\mathbf{g}^{i}$ is used to capture features that are not assigned to any prototypical part. Importantly, these features do not necessarily correspond to background pixels. Instead, they include informative regions of the object that are not consistently detected as one of the predefined parts by the part-detection module. As illustrated in Fig. 12, these non-prototypical concepts often correspond to distinctive yet less frequent features that are still valuable for decision-making. For example:
  • In the Cars dataset, the “cargo” region is activated as a non-prototypical concept. Though it appears in a small subset of vehicles only, that provides strong discriminative power.
  • In the CUB dataset, the “webbed feet” of certain birds (e.g., Albatross species) are highlighted. These regions are highly specific and explainable, yet do not form a consistent enough prototype across classes to be captured in the prototypical set.
@@ C Acc F(1)-F(5) C Acc F(1)-F(5)
Table 5: Impact of prototypical ($\mathbf{z^{i}}$) and non-prototypical ($\mathbf{g^{i}}$) concept vectors on PCMNet’s classification accuracy (C Acc) and faithfulness scores (F(1)-F(5)) on Cars and Birds datasets. The full model achieves superior performance, highlighting the benefit of combining both concept types.
Comparison of Performance, Interpretability, and Efficiency
To better visualize the trade-offs among classification performance, interpretability, and computational efficiency, we present a summary plot comparing PCMNet to recent interpretable AI baselines (Fig. 9). Each method is represented using a unique shape, and the color intensity of each point reflects its FLOPs (darker points indicate higher computational cost). The x-axis corresponds to classification accuracy on the Stanford Cars dataset, while the y-axis shows faithfulness under occlusion (F(3)), one of our main interpretability metrics.
Figure 9: Trade-off comparison between accuracy, faithfulness (F(3)), and efficiency (FLOPs) across methods on the Stanford Cars dataset. Each method is represented by a distinct marker shape, and the color intensity encodes its computational cost (darker = less FLOPs). PCMNet achieves the best balance of high accuracy, top interpretability, and low computational cost, demonstrating its practical advantage over state-of-the-art interpretable models.
PCMNet (full) achieves an excellent balance: it attains competitive accuracy (90.2%) and the highest interpretability (F(3) = 67.4) among all methods, while maintaining one of the lowest FLOPs (17.4G). In contrast, ProtoViT achieves slightly higher accuracy (91.8%) but suffers from lower interpretability (F(3) = 44.6) and significantly higher FLOPs (24.8G). ProtoPNet and PiPNet either underperform in accuracy or require substantially more computation. This visualization supports our claim that PCMNet is highly practical: it delivers state-of-the-art interpretability without sacrificing classification performance or efficiency.
Appendix C: Additional Qualitative Results
More results of the PCMNet prototypical concept are shown in Figs. 10 and 11. Fig 12 illustrates examples of non-prototypical concepts discovered by our model for the Cars (left) and CUBs (right) datasets. For each dataset, the concept location on a test image is highlighted with a green bounding box, while corresponding activations in training examples are shown on the right using red bounding boxes. These concepts provide complementary cues to prototypical ones. For example, in the first row of Cars, the ”cargo” attributes are not well represented in prototypical concepts since they appear in a small subset of vehicles, yet they are highly discriminative for decision-making. Similarly, the last row in CUBs shows the ”webbed feet” concept, which is especially relevant and explainable for Albatross species, while foot prototypes are not sufficiently distinctive across the dataset. These findings highlight the value of incorporating non-prototypical concepts in improving both interpretability and robustness of the model.
Figure 10: Visualization of activated concepts in test data and showing some examples in the training set for StanfordCars datasets. Each concept is shown with an index and percentage of explanation for the model decision.
Figure 11: Visualization of activated concepts in test data and showing some examples in the training set for CUB-200 Birds datasets. Each concept is shown with an index and percentage of explanation for the model decision.
Figure 12: Visualization of non-prototypical activated concepts in test data and showing some examples in the training set for (left) Stanford Cars and (right) CUB-200 Birds datasets. Similar to ProtoPNet and PIPNet, we select the patch for the region of the activated concept.